# Every finite set of natural numbers is realizable as algebraic periods of a Morse–Smale diffeomorphism

### v1

@@ 4.2. An estimate of the number of algebraically finite type classes
More information about the functions $P(N)$ one can find in [Wei1]. At this moment let us only mention the asymptotic behavior of it (Hardy-Ramanujan 1918):
@@
The number of algebraically finite type mapping classes of an orientable or non-orientable closed surface of genus $\mathrm{g}$ is estimated from below by $P(\mathrm{g})$, the number of unrestricted partitions of $\mathrm{g}$. Consequently, there are at least $P(\mathrm{g})$ homotopy classes of Morse–Smale diffeomorphisms on $S_{\mathrm{g}}$ and on $N_{\mathrm{g}}$.
@@
where $a_{n}(f)$ denote the coefficients in the $k$-periodic expansion of $f$.
We apply the construction from the proof of Theorem 3.1. Recall that in the orientation-preserving case (1) we realized ${\mathcal{A}}$ as ${\mathcal{A}P}(f)$ of a map $f\colon S_{\mathrm{g}}\to S_{\mathrm{g}}$ such that $a_{k}(f)=-2$ for $k\in{\mathcal{A}}\setminus\{1\}$, $a_{1}(f)=2$ if $1\in{\mathcal{A}}$ and obviously $a_{k}(f)=0$ for $k\notin{\mathcal{A}}$. From the description of the genus or just from the construction it follows that $2\mathrm{g}=2-\sum_{k}a_{k}(f)\cdot k$ and so one can easily check that $\mathrm{g}=\sum_{k}p_{k}k$ by the formula (10). Recall that the surface $S_{\mathrm{g}}$ was formed from the pieces $\Sigma_{k}$ for $k\in{\mathcal{A}}^{\prime}$ and connecting cylinders.
Now, for a given partition $(p_{1},\ldots,p_{\mathrm{g}})$ of $\mathrm{g}$ and every $k$ we take $p_{k}$ copies of the surface $\Sigma_{k}$ of genus $k$ together with defined $k$-periodic homeomorphisms and glue them appropriately as in Theorem 3.1. From the construction it is straightforward that $\mathrm{g}=\sum p_{k}k$ is the genus of the resulted surface and $\mathrm{t}rH_{1}(f^{n})=\sum_{k}2p_{k}{\rm reg}_{k}(n).$ Therefore we obtained an algebraically finite type homeomorphism $f$ of $S_{\mathrm{g}}$ such that
@@
is consistent with the formula (10).
Now consider the non-orientble case. The analogous correspondence between partitions $(p_{1},\ldots,p_{\mathrm{g}})$ of $\mathrm{g}$ and homeomorphisms $f\colon N_{\mathrm{g}}\to N_{\mathrm{g}}$ of algebraically finite type is given by
@@
Recall that in the construction in the proof of Theorem 3.1 in the non-orientable case (3) we realized ${\mathcal{A}}$ as ${\mathcal{A}P}(f)$ of $f\colon N_{\mathrm{g}}\to N_{\mathrm{g}}$ such that $a_{k}(f)=-1$ for $k\in{\mathcal{A}}\setminus\{1\}$. Similarly, for a given partition $(p_{1},\ldots,p_{\mathrm{g}})$ of $\mathrm{g}$ we repeat the construction with $p_{k}$ copies of the non-orientable surface $\Sigma_{k}$ of genus $k$ on which the obtained function is $k$-periodic. Therefore the obtained surface is clearly of genus $\mathrm{g}=\sum p_{k}k$ and by the computations during the proof of Theorem 3.1 we get
@@
Thus $f$ is an algebraically finite type homeomorphism of $N_{\mathrm{g}}$ such that
@@
where the coefficients $a_{k}(f)$ of the $k$-periodic expansion of $f$ are determined by (11).
In both the cases, by Theorem 2.3 of da Rocha the mapping class of $f$ contains a Morse–Smale representative. The relations (10) and (11) show that diffeomorphisms corresponding in our construction to different unrestricted partitions of $\mathrm{g}$ have different $k$-periodic expansions, and consequently are not homotopic, i.e. they are in different mapping classes. ∎
5. Appendix - Algebraic periods and minimal Lefschetz periods
We remind that the zeta function for the sequence of Lefschetz numbers $L=\{l_{n}\}_{n}:=\{L(f^{n})\}_{n}$ of iterations is defined as:
@@ 5. Appendix - Algebraic periods and minimal Lefschetz periods
In [Ll-Si1] the authors defined the set of minimal Lefschetz periods of a diffeomorphism $f\colon M\to M$, denoted ${\rm MPer}_{L}(f)$, in the following way.
Definition 5.1.
@@
where $\Delta_{i}=\pm 1$, i.e. we take into account rational representations of $\zeta(L;z)$ as products of powers and polynomials $(1+\Delta_{i}\,z^{r_{i}})$.
Geometrically, it known (see generalized Franks formula being a consequence of (3.3.9) in [Je-Ma]) that $\zeta(L;z)$ has such a representation, with each term related to some periodic orbit in case of transversal maps.
However, the factors and its amount are not canonically determined by the sequence $\{L(f^{n})\}_{n}$ in this formula. Llibre and coauthors derived the set ${\rm MPer}_{L}(f)$ for several examples (see the references of [Ll-Si3]) observing meantime that it does not contain even numbers by a simple algebraic argument included here in the proof of Proposition 5.2 (cf. [Ll-Si2]). Notify also, that formally we have infinitely many of such rational representations potentially with several possible values $\eta$, $r_{i}$, and $m_{i}$, which geometrically reflects the fact that it may appear periodic orbits of high periods whose contributions to $\zeta(L,z)$ annihilate one another.
Recently, in [GLM] the authors showed that for a transversal map $f\in C_{T}(M)$ of a manifold $M$ we have ${\rm MPer}_{L}(f)={\mathcal{A}P}(f)\cap\{2\mathbb{N}-1\}={\mathcal{A}P}_{odd}(f)$ using the $k$-periodic expansion of the sequence $\{L(f^{n})\}$. Originally this theorem is stated for Morse–Smale diffeomorphisms, but the argument holds for the class of transversal maps. It is worth pointing out that in [GLM] is used the fact that all summands in representation of $\{L(f^{n})\}_{n}$ come from the geometric representation of $\zeta(L,z)$. Below we present another purely algebraic way of showing that ${\rm MPer}_{L}(f)={\mathcal{A}P}_{odd}(f)$ for a larger class of maps.
By the definition, the Lefschetz zeta function $\zeta(L;z)\in\mathbb{Z}[[z]]$ is a power series which is a rational function over $\mathbb{Z}$, i.e. an element of $\mathbb{Z}(z)$ the field of fractions of the ring of integral polynomials, as follows from the homology formula for the Lefschetz numbers (cf. [Je-Ma, (3.1.27)]). Moreover, $\zeta(L;z)$ has the following multiplicative representation (cf. [Je-Ma, (3.1.22)])
@@
Note that in the case when the sequence $\{L(f^{n})\}_{n}$ is bounded, e.g. if $f$ is a Morse–Smale diffeomorphism, the set ${\mathcal{A}P}(f)=\{n:a_{n}(f)\neq 0\}$ is finite, and consequently the product (14) is finite.
Moreover, the formula (14) is a unique rational expression of $\zeta_{f}(z)$ in terms of polynomials $(1-z^{n})$ since none of the polynomials $(1-z^{n})$ can be expressed as a rational function of polynomials $(1-z^{k})$ for $k\neq n$. Indeed, let $(1-z^{n})=\prod_{k\neq n}(1-z^{k})^{c_{k}}$ for $c_{k}\in{\mathbb{Z}}$ and finitely many $c_{k}\neq 0$, and let $k_{0}$ be the largest number such that $c_{k_{0}}\neq 0$.
  • if $k_{0}<n$, then the left-hand side has a zero at a primitive root of unity of degree $n$, but the right-hand side does not,
@@
Proposition 5.2.
Let $f\colon X\to X$ be a map of a finite CW-complex $X$ such that the sequence $\{L(f^{n})\}_{n}$ of Lefschetz numbers of iterations is bounded. Then
@@
    where $e_{k}=c_{k}+d_{k/2}-d_{k}$ if $k$ is even and $e_{k}=c_{k}-d_{k}$ if $k$ is odd. Since the form (14) is unique, $e_{k}=a_{k}(f)$. This means that $c_{n}-d_{n}=a_{n}(f)\neq 0$ for $n\in{\mathcal{A}P}_{odd}(f)$. In particular $c_{n}\neq 0$ or $d_{n}\neq 0$, so $n\in{\rm MPer}_{L}(f)$ because (15) is arbitrary.

### v2

@@ 4.2. An estimate of the number of conjugacy classes of algebraically finite type mapping classes
More information about the functions $P(N)$ can be found in [28]. At this moment, let us only mention the asymptotic behavior of it (Hardy-Ramanujan 1918):
@@
The number of conjugacy classes of algebraically finite type mapping classes of an orientable or non-orientable closed surface of genus $\mathrm{g}$ is estimated from below by $P(\mathrm{g})$, the number of unrestricted partitions of $\mathrm{g}$. Consequently, there are at least $P(\mathrm{g})$ conjugacy classes of ($S_{\mathrm{g}}$ and $N_{\mathrm{g}}$) mapping classes containing Morse–Smale diffeomorphisms.
@@
where $a_{n}(f)$ denote the coefficients in the periodic expansion of $f$.
We apply the construction from the proof of Theorem 3.1. Recall that in the orientation-preserving case (1), we realized ${\mathcal{A}}$ as ${\mathcal{A}P}(f)$ of a map $f\colon S_{\mathrm{g}}\to S_{\mathrm{g}}$ such that $a_{k}(f)=-2$ for $k\in{\mathcal{A}}\setminus\{1\}$, $a_{1}(f)=2$ if $1\in{\mathcal{A}}$, and obviously $a_{k}(f)=0$ for $k\notin{\mathcal{A}}$. From the description of the genus, or just from the construction, it follows that $2\mathrm{g}=2-\sum_{k}a_{k}(f)\cdot k$, and so one can easily check that $\mathrm{g}=\sum_{k}p_{k}k$ by formula (10). Recall that the surface $S_{\mathrm{g}}$ was formed from the pieces $\Sigma_{k}$ for $k\in{\mathcal{A}}^{\prime}$ and connecting cylinders.
Now, for a given partition $(p_{1},\ldots,p_{\mathrm{g}})$ of $\mathrm{g}$ and every $k$, we take $p_{k}$ copies of the surface $\Sigma_{k}$ of genus $k$ together with defined periodic homeomorphisms, and glue them appropriately as in Theorem 3.1. From the construction, it is straightforward that $\mathrm{g}=\sum p_{k}k$ is the genus of the resulted surface, and $\mathrm{t}rH_{1}(f^{n})=\sum_{k}2p_{k}{\rm reg}_{k}(n).$ Therefore, we obtained an algebraically finite type homeomorphism $f$ of $S_{\mathrm{g}}$ such that
@@
is consistent with formula (10).
Now, consider the non-orientble case. The analogous correspondence between partitions $(p_{1},\ldots,p_{\mathrm{g}})$ of $\mathrm{g}$ and homeomorphisms $f\colon N_{\mathrm{g}}\to N_{\mathrm{g}}$ of algebraically finite type is given by
@@
Recall that in the construction in the proof of Theorem 3.1 in the non-orientable case (3), we realized ${\mathcal{A}}$ as ${\mathcal{A}P}(f)$ of $f\colon N_{\mathrm{g}}\to N_{\mathrm{g}}$ such that $a_{k}(f)=-1$ for $k\in{\mathcal{A}}\setminus\{1\}$. Similarly, for a given partition $(p_{1},\ldots,p_{\mathrm{g}})$ of $\mathrm{g}$, we repeat the construction with $p_{k}$ copies of the non-orientable surface $\Sigma_{k}$ of genus $k$ on which the obtained function is periodic. Therefore, the obtained surface is clearly of genus $\mathrm{g}=\sum p_{k}k$, and by the computations during the proof of Theorem 3.1, we get
@@
Thus, $f$ is an algebraically finite type homeomorphism of $N_{\mathrm{g}}$ such that
@@
where the coefficients $a_{k}(f)$ of the periodic expansion of $f$ are determined by (11).
In both cases, by Theorem 2.3 of da Rocha, the mapping class of $f$ contains a Morse–Smale representative. The relations (10) and (11) show that diffeomorphisms corresponding in our construction to different unrestricted partitions of $\mathrm{g}$ have different periodic expansions, and consequently their homotopy classes are not conjugated. ∎
Appendix A Algebraic periods and minimal sets of Lefschetz periods
We recall that the zeta function for the sequence of Lefschetz numbers $L=(l_{n}):=(L(f^{n}))$ of iterations is defined as
@@ Appendix A Algebraic periods and minimal sets of Lefschetz periods
In [16], the authors defined the minimal set of Lefschetz periods of a diffeomorphism $f\colon M\to M$, denoted ${\rm MPer}_{L}(f)$, in the following way.
Definition A.1.
@@
where $r_{i},r_{j}^{\prime},r_{k}^{\prime\prime}\in{\mathbb{N}}$, $m_{i},\,m_{j}^{\prime},m_{k}^{\prime\prime}\in{\mathbb{Z}}$, and $\Delta_{i}=\pm 1$, i.e. we take into account rational representations of $\zeta(L;z)$ as products of powers of polynomials $(1+\Delta_{i}\,z^{r_{i}})$.
Geometrically, it is known (see the generalized Franks formula being a consequence of (3.3.9) in [14]) that $\zeta(L;z)$ has such a representation, with each term related to some periodic orbit in case of transversal maps.
However, the factors and their amount are not canonically determined by the sequence $(L(f^{n}))$ in this formula. Llibre and coauthors derived the set ${\rm MPer}_{L}(f)$ for several examples (see the references of [18]), observing that it does not contain even numbers by a simple algebraic argument included here in the proof of Proposition A.2 (cf. [17]). Note also that, formally, we have infinitely many of such rational representations, potentially with several possible values $\eta$, $r_{i}$, and $m_{i}$, which geometrically reflects the fact that periodic orbits of high periods may appear, whose contributions to $\zeta(L,z)$ annihilate one another.
Recently, in [10], the authors showed that for a transversal map $f\in C_{T}(M)$ of a manifold $M$, we have ${\rm MPer}_{L}(f)={\mathcal{A}P}(f)\cap(2{\mathbb{N}}-1)={\mathcal{A}P}_{odd}(f)$ using the periodic expansion of the sequence $(L(f^{n}))$. Originally, this theorem is stated for Morse–Smale diffeomorphisms, but the argument holds for the class of transversal maps. It is worth pointing out that in [10] the fact that all summands in representation of $(L(f^{n}))$ come from the geometric representation of $\zeta(L,z)$ is used. Below, we present another purely algebraic way of showing that ${\rm MPer}_{L}(f)={\mathcal{A}P}_{odd}(f)$ for a larger class of maps.
The Lefschetz zeta function $\zeta(L;z)$ is a rational function over ${\mathbb{Z}}$ (cf. [14, (3.1.27)]). Moreover, $\zeta(L;z)$ has the following multiplicative representation (cf. [14, (3.1.22)]):
@@
Note that in the case when the sequence $(L(f^{n}))$ is bounded, e.g. if $f$ is a Morse–Smale diffeomorphism, the set ${\mathcal{A}P}(f)=\{n:a_{n}(f)\neq 0\}$ is finite, and consequently the product (14) is finite.
Moreover, the formula (14) is a unique rational expression of $\zeta_{f}(z)$ in terms of polynomials $(1-z^{n})$ since none of the polynomials $(1-z^{n})$ can be expressed as a rational function of polynomials $(1-z^{k})$ for $k\neq n$. Indeed, let $(1-z^{n})=\prod_{k\neq n}(1-z^{k})^{c_{k}}$ for $c_{k}\in{\mathbb{Z}}$, and finitely many $c_{k}\neq 0$, and let $k_{0}$ be the largest number such that $c_{k_{0}}\neq 0$.
  • If $k_{0}<n$, then the left-hand side has a zero at a primitive root of unity of degree $n$, but the right-hand side does not.
@@
Proposition A.2.
Let $f\colon X\to X$ be a map of a finite CW-complex $X$ such that the sequence $(L(f^{n}))$ of Lefschetz numbers of iterations is bounded. Then,
@@
    where $e_{k}=c_{k}+d_{k/2}-d_{k}$ if $k$ is even and $e_{k}=c_{k}-d_{k}$ if $k$ is odd. Since the form (14) is unique, $e_{k}=a_{k}(f)$. This means that $c_{n}-d_{n}=a_{n}(f)\neq 0$ for $n\in{\mathcal{A}P}_{odd}(f)$. In particular, $c_{n}\neq 0$ or $d_{n}\neq 0$, so $n\in{\rm MPer}_{L}(f)$ because (15) is arbitrary.